# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

### v4

@@ Abstract
Long-term, high-fidelity simulation of slow-changing physical systems, such as the ocean and climate, presents a fundamental challenge in scientific computing. Traditional autoregressive machine learning models often fail in these tasks as minor errors accumulate and lead to rapid forecast degradation. To address this problem, we propose NeuralOM, a general neural operator framework designed for simulating complex, slow-changing dynamics. NeuralOM’s core consists of two key innovations: (1) a Progressive Residual Correction Framework that decomposes the forecasting task into a series of fine-grained refinement steps, effectively suppressing long-term error accumulation; and (2) a Physics-Guided Graph Network whose built-in adaptive messaging mechanism explicitly models multi-scale physical interactions, such as gradient-driven flows and multiplicative couplings, thereby enhancing physical consistency while maintaining computational efficiency. We validate NeuralOM on the challenging task of global Subseasonal-to-Seasonal (S2S) ocean simulation. Extensive experiments demonstrate that NeuralOM not only surpasses state-of-the-art models in forecast accuracy and long-term stability, but also excels in simulating extreme events. For instance, at a 60-day lead time, NeuralOM achieves a 13.3% lower RMSE compared to the best-performing baseline, offering a stable, efficient, and physically-aware paradigm for data-driven scientific computing. Code link: https://github.com/YuanGao-YG/NeuralOM.
@@ Introduction
The long-term simulation of complex physical systems (wang2024modeling; wu2025advanced), such as the global ocean and climate, is a cornerstone of modern science yet remains a great challenge. While machine learning (ML) models offer a computationally efficient path to complement traditional numerical solvers, they face a fundamental bottleneck in such tasks: long-term stability. This issue is especially severe in slow-changing systems, where the state evolves in a subtle manner. In standard autoregressive frameworks, minor prediction errors at each step inevitably accumulate, causing the simulation to diverge catastrophically from the true physical dynamics, a well-known problem of compounding error. This error cascade often leads data-driven models to fail in long-term rollouts. As shown in Figure 1 (bottom left), a state-of-the-art baseline model generates substantial, physically unrealistic artifacts in its long-range forecast.
@@
To address this challenge, existing machine learning approaches have been explored, but they are limited in two key aspects. ❶. First, they often struggle to capture the fine-grained dynamics inherent to slow-changing systems. Many models are designed to predict sharp, significant shifts, making them ill-suited for learning the small, incremental updates that define ocean dynamics. This typically results in either over-smoothing the predictions or failing to capture critical physical anomalies. ❷. Second, the internal mechanisms of these models are often physics-agnostic. For instance, standard message passing in Graph Neural Networks (GNNs) (keisler2022forecasting) treats all node interactions uniformly, unable to distinguish between fundamentally different physical processes like gradient-driven flows and multiplicative couplings between variables. Furthermore, they lack the ability to adaptively model phenomena across different spatio-temporal scales, from large-scale, slow-moving currents to small-scale, energetic eddies.
@@
Experimental Observations. ❶ State-of-the-art accuracy in long-term simulation: As shown in Table 1, in $60$-day simulation, NeuralOM achieves a normalized RMSE of just $0.7015$, a 13.313.3% reduction compared to the best-performing baseline, WenHai ($0.8093$). It also improves the Anomaly Correlation Coefficient (ACC) by 12.012.0%, indicating more accurate tracking of the physical field’s evolution. ❷ Superior visual fidelity and physical consistency: As illustrated in Figure 1 and Figure 3, unlike baseline models like CirT that collapse due to error explosion, NeuralOM produces physically consistent forecasts that remain highly aligned with the ground truth even at $60$ days, successfully preserving critical fine-scale structures. ❸ Validated effectiveness of our designs through ablation studies: As detailed in Table 3 and Table 4, removing any of our core components, either the Progressive Residual Correction Framework or the Physics-Guided Adaptive Graph Messaging, leads to a significant drop in performance. In particular, incorporating climatology priors improves the ACC metric by more than $5$-fold, highlighting that every design choice is crucial to achieve accurate and stable simulations.
@@
In recent years, previous works in spatio-temporal mining have achieved stupendous achievements (pfaff2020learning; li2020fourier; wu2023earthfarseer; wu2024pastnet; wang2024nuwadynamics; wu2024prometheus; wu2024pure; wu2025turb; wu2025spatiotemporal; wu2025differential; jia2025learning). However, many of these methods are only applicable in idealized or relative simple scenario, whereas ocean system is a complex dynamical system in the real world, requiring the design of specialized methods for accurate prediction (wu2024neural). Research in data-driven ocean simulation is primarily divided into two streams: task-specific predictive models and general-purpose global foundation models.
@@
These models aim to simulate specific oceanic or atmospheric phenomena and have achieved impressive results in their respective domains. Examples include work on El Nio-Southern Oscillation (ENSO) prediction (ham2019deep; hu2021deep; chang2023predicting; lyu2024resonet; chen2025combined), as well as models for the Madden-Julian Oscillation (MJO) (kim2021deep; delaunay2022interpretable; yang2023exploring; shin2024data; shin2024deep) and Marine Heatwaves (MHWs) (jacox2022global; sun2024deep; parasyris2025marine). The success of these models validates the immense potential of machine learning for complex geoscience problems. However, their limitation is also apparent: they are typically highly optimized for a single objective and often overlook the complex interactions between different physical variables and across multiple scales. This makes them difficult to generalize into a foundation model capable of capturing the full dynamics of the ocean system.
@@
Inspired by the success of foundation models in atmospheric science (pathak2022FourCastNet; bi2023accurate; chen2023fengwu; chen2023fuxi; lam2023learning; oskarsson2024probabilistic; nguyen2024scaling; kochkov2024neural; gao2025oneforecast), a series of global ocean foundation models have recently emerged, such as AI-GOMS (xiong2023ai), Xihe (wang2024xihe), LangYa (yang2024langya), and WenHai (cui2025forecasting). These models aim to build a unified digital twin of the ocean and have already surpassed traditional numerical methods in computational efficiency and, in some cases, accuracy (hao2025deep). However, they still face the two core challenges we identified in the introduction. First, they generally adopt standard autoregressive frameworks and have not fundamentally solved the problem of error accumulation in long-term simulations, leading to a decline in forecast fidelity over time. Second, their designs often lack sufficient physical constraints. For example, while models like WenHai (cui2025forecasting) incorporate domain knowledge into a Swin Transformer (liu2021swin), their feature interaction mechanisms remain relatively physics-agnostic, failing to explicitly model physical processes like gradient changes or variable couplings, nor can they adaptively handle multi-scale dynamics.
@@
$\hat{\mathbf{O}}_{t+1}^{(1)}$ $=\mathcal{M}_{\text{base}}(\mathbf{X}_{t})$ (2)
@@
We train each component of NeuralOM by minimizing the relative mean squared error (MSE) between the predicted and ground-truth states. The loss function $\mathcal{L}$ is defined as:
$\mathcal{L}(\Theta)=\mathbb{E}_{t,i,j,k}\left[\frac{\left(\hat{O}_{i,j,k}^{t+1}-O_{i,j,k}^{t+1}\right)^{2}}{\mathbb{E}_{i,j,k}[(O_{i,j,k}^{t+1})^{2}]+\epsilon}\right],$ (6)
where $\hat{O}$ and $O$ are the predicted and ground-truth values, respectively, indexed over time $t$, spatial locations $(i,j)$, and variable channels $k$. The expectation in the denominator serves as a normalization constant to balance the influence of variables with different magnitudes, and $\epsilon$ is a small constant for numerical stability.
@@
For long-range simulation and forecasting, we employ a standard autoregressive (or “rollout”) strategy. Starting from an initial state $\mathbf{X}_{0}$, the model continuously feeds its own previous prediction as input for the next step to generate a continuous trajectory spanning several weeks or months. For idealized simulation experiments, we use ground-truth atmospheric data as forcing. This approach eliminates interference from atmospheric forecast errors, allowing for a fairer assessment of the ocean models’ intrinsic performance. For more realistic forecasting experiments, we drive NeuralOM with forecast fields generated by a state-of-the-art atmospheric forecasting model, OneForecast (gao2025oneforecast).
Summary.
@@
FourCastNet (pathak2022FourCastNet) arXiv 2022 5.6336 0.0781 >10>10 0.0189 >10>10 0.0046 >10>10 0.0007
CirT (liu2025cirt) ICLR 2025 2.7525 0.0143 2.8700 0.0115 2.9528 0.0099 3.0645 0.0081
WenHai (cui2025forecasting) Nature Communications 2025 0.7200 0.5067 0.7468 0.4705 0.7702 0.4386 0.8093 0.3852
NeuralOM (ours) 0.6510 0.5299 0.6666 0.5010 0.6797 0.4756 0.7015 0.4316
NeuralOM Improvement 9.59% 4.58% 10.74% 6.49% 11.74% 8.43% 13.31% 12.02%
@@
Our experiments are based on the GLORYS12 global ocean reanalysis dataset, which provides daily mean ocean states from 1993 to 2023. We split the data into a training set (1993-2017), a validation set (2018-2019), and a test set (2020). The core task is autoregressive forecasting at the Subseasonal-to-Seasonal (S2S) scale, with rollout horizons ranging from several days up to 60 days. The input consists of 93 ocean state variables (including salinity, velocity, temperature at various depths, and sea surface height) and 8 atmospheric variables as external forcing. All spatial data is downsampled to a 24-hour temporal resolution (daily mean for ocean and 12:00 UTC for forcings) and 0.5° spatial resolution. Further details on data processing and variables are available in the Appendix.
@@
We compare NeuralOM against three representative models: (1) FourCastNet (pathak2022FourCastNet), a widely recognized atmospheric science foundation model based on Fourier Neural Operators. (2) CirT (liu2025cirt), a state-of-the-art atmospheric forecasting model designed for the S2S scale. (3) WenHai (cui2025forecasting), a state-of-the-art ML-based ocean model that has demonstrated superior performance over traditional numerical models and ML-based ocean foundation models. To ensure a fair comparison, all models, except for the officially released WenHai results, were retrained under our unified framework.
@@
We use two primary metrics to assess performance: (1) Root Mean Square Error (RMSE) to measure the average prediction error, and (2) Anomaly Correlation Coefficient (ACC) to evaluate the model’s ability to capture deviations from the climatological mean (i.e., anomalies). We report the normalized RMSE and ACC, averaged over all 93 ocean variables. Additionally, we use the Critical Success Index (CSI) and the Symmetric Extremal Dependence Index (SEDI) to evaluate the model’s performance on extreme events.
@@
To assess practical applicability, we drive the ocean models with atmospheric forecasts from OneForecast, which introduces real-world uncertainty. As common atmospheric forecast models don’t provide all forcings required by WenHai at the S2S scale, we can’t assess its results. Even in this more challenging scenario, Table 2 shows that NeuralOM retains its lead, achieving significantly better RMSE ($0.68$) and ACC ($0.49$) at a $35$-day lead time. This demonstrates the robustness of NeuralOM for real-world applications, further supported by the high-fidelity visuals in Figure 5(a).
@@
FourCastNet 0.56 0.68 0.69 0.47 1.54 0.16
CirT 1.61 0.11 2.17 0.04 2.55 0.02
Ours 0.53 0.73 0.62 0.60 0.68 0.49
@@
Variants RMSE ACC
NeuralOM w/o PRC 0.6692 0.5503
NeuralOM w/o PEI 0.6865 0.5168
NeuralOM w/o MANA Sum 0.7037 0.5361
NeuralOM w/o MANA Mean 0.6978 0.5161
NeuralOM 0.6097 0.5983
NeuralOM excels at capturing high-impact extreme events.
VARIANTS 23 Layers S 23 Layers T SSH
RMSE ACC RMSE ACC RMSE ACC
NeuralOM w/o Subtract Climatology 0.0515 0.1061 0.1518 0.4535 0.2424 0.5747
NeuralOM W Subtract Climatology 0.0105 0.6412 0.0985 0.6508 0.1374 0.6695
@@
To deeply understand the driving factors behind NeuralOM’s superior performance and to validate the effectiveness of each of our design choices, we conduct a series of detailed ablation studies. We primarily investigate two aspects: the importance of the input representation (i.e., subtracting climatology) and the contribution of each component in our core architecture.
@@
As mentioned in our methodology, we argue that focusing the model on learning “anomalies” rather than the full state is crucial for capturing slow-changing dynamics. To verify this, we train a variant of NeuralOM without subtracting the climatological mean. The results, shown in Table 4, are striking. For sea salinity (S) across 23 depth layers, the model with climatology priors (W Subtract Climatology) achieved an ACC of 0.6412, whereas the model without them (W/O Subtract Climatology) only reached 0.1061—a performance gap of over 5-fold. This result provides strong evidence that incorporating physical priors like climatology into the input representation is a prerequisite for enabling the model to effectively learn fine-grained physical signals and achieve high-fidelity simulations.
@@
Next, we evaluate the individual contributions of NeuralOM’s core components by removing them one by one. The results, summarized in Table 3, show that removing any single component leads to a significant degradation in performance:
Without Progressive Residual Correction (w/o PRC): When we replaced the multi-stage PRC framework with a single, deeper model, the RMSE increased from 0.6097 to 0.6692. This demonstrates that decomposing the complex prediction task into a series of simpler residual correction steps is crucial for suppressing error accumulation and ensuring long-term stability.
Without Physics-inspired Edge-Interactions (w/o PEI): Removing our physics-inspired interaction module (PEI), thereby reverting to a traditional, physics-agnostic message passing scheme, caused the ACC to drop from 0.5983 to 0.5168. This directly validates that explicitly modeling physical processes like gradients and couplings in message passing is effective for enhancing the model’s physical consistency.
Multi-scale Adaptive Node Aggregation (w/o MANA Sum/Mean): Removing either the sum or the mean aggregation path individually resulted in degraded performance. This confirms the necessity of our adaptive aggregation mechanism, which allows the model to flexibly handle ocean dynamics at different scales.
In summary, these ablation studies clearly demonstrate that the success of NeuralOM is not coincidental but is the result of a synergistic combination of its carefully designed components from the input representation and macroscopic framework to the microscopic interaction mechanisms.

### v5

@@ Abstract
Long-term, high-fidelity simulation of slow-changing physical systems, such as the ocean and climate, presents a fundamental challenge in scientific computing. Traditional autoregressive machine learning models often fail in these tasks as minor errors accumulate and lead to rapid forecast degradation. To address this problem, we propose NeuralOM, a general neural operator framework designed for simulating complex, slow-changing dynamics. NeuralOM’s core consists of two key innovations: (1) a Progressive Residual Correction Framework that decomposes the forecasting task into a series of fine-grained refinement steps, effectively suppressing long-term error accumulation; and (2) a Physics-Guided Graph Network whose built-in adaptive messaging mechanism explicitly models multi-scale physical interactions, such as gradient-driven flows and multiplicative couplings, thereby enhancing physical consistency while maintaining computational efficiency. We validate NeuralOM on the challenging task of global Subseasonal-to-Seasonal (S2S) ocean simulation. Extensive experiments demonstrate that NeuralOM not only surpasses state-of-the-art models in forecast accuracy and long-term stability, but also excels in simulating extreme events. For instance, at a 60-day lead time, NeuralOM achieves a 13.3% lower RMSE compared to the best-performing baseline, offering a stable, efficient, and physically-aware paradigm for data-driven scientific computing.
Codehttps://github.com/YuanGao-YG/NeuralOM
@@ Introduction
The long-term simulation of complex physical systems (Wang et al. 2024b; Wu et al. 2025c), such as the global ocean and climate, is a cornerstone of modern science yet remains a great challenge. While machine learning (ML) models offer a computationally efficient path to complement traditional numerical solvers, they face a fundamental bottleneck in such tasks: long-term stability. This issue is especially severe in slow-changing systems, where the state evolves in a subtle manner. In standard autoregressive frameworks, minor prediction errors at each step inevitably accumulate, causing the simulation to diverge catastrophically from the true physical dynamics, a well-known problem of compounding error. This error cascade often leads data-driven models to fail in long-term rollouts. As shown in Figure 1 (bottom left), a state-of-the-art baseline model generates substantial, physically unrealistic artifacts in its long-range forecast.
@@
To address this challenge, existing machine learning approaches have been explored, but they are limited in two key aspects. ❶. First, they often struggle to capture the fine-grained dynamics inherent to slow-changing systems. Many models are designed to predict sharp, significant shifts, making them ill-suited for learning the small, incremental updates that define ocean dynamics. This typically results in either over-smoothing the predictions or failing to capture critical physical anomalies. ❷. Second, the internal mechanisms of these models are often physics-agnostic. For instance, standard message passing in Graph Neural Networks (GNNs) (Keisler 2022) treats all node interactions uniformly, unable to distinguish between fundamentally different physical processes like gradient-driven flows and multiplicative couplings between variables. Furthermore, they lack the ability to adaptively model phenomena across different spatio-temporal scales, from large-scale, slow-moving currents to small-scale, energetic eddies.
@@
Experimental Observations. ❶ State-of-the-art accuracy in long-term simulation: As shown in Table 1, in $60$-day simulation, NeuralOM achieves a normalized RMSE of just $0.7014$, a 13.313.3% reduction compared to the best-performing baseline, WenHai ($0.8091$). It also improves the Anomaly Correlation Coefficient (ACC) by 12.012.0%, indicating more accurate tracking of the physical field’s evolution. ❷ Superior visual fidelity and physical consistency: As illustrated in Figure 1 and Figure 3, unlike baseline models like CirT that collapse due to error explosion, NeuralOM produces physically consistent forecasts that remain highly aligned with the ground truth even at $60$ days, successfully preserving critical fine-scale structures. ❸ Validated effectiveness of our designs through ablation studies: As detailed in Table 2 and Table 3, removing any of our core components, either the Progressive Residual Correction Framework or the Physics-Guided Adaptive Graph Messaging, leads to a significant drop in performance. In particular, incorporating climatology priors improves the ACC metric by more than $5$-fold, highlighting that every design choice is crucial to achieve accurate and stable simulations.
@@
In recent years, previous works in spatio-temporal mining have achieved stupendous achievements (Pfaff et al. 2020; Li et al. 2020; Wu et al. 2023, 2024d; Wang et al. 2024a; Wu et al. 2024b, a, 2025b, 2025a, 2025d; Jia et al. 2025). However, many of these methods are only applicable in idealized or relative simple scenario, whereas ocean system is a complex dynamical system in the real world, requiring the design of specialized methods for accurate prediction (Wu et al. 2024c). Research in data-driven ocean simulation is primarily divided into two streams: task-specific predictive models and general-purpose global foundation models.
@@
These models aim to simulate specific oceanic or atmospheric phenomena and have achieved impressive results in their respective domains. Examples include work on El Nio-Southern Oscillation (ENSO) prediction (Ham et al. 2019; Hu et al. 2021; Chang et al. 2023; Lyu et al. 2024; Chen et al. 2025), as well as models for the Madden-Julian Oscillation (MJO) (Kim et al. 2021; Delaunay and Christensen 2022; Yang et al. 2023; Shin et al. 2024a, b) and Marine Heatwaves (MHWs) (Jacox et al. 2022; Sun et al. 2024; Parasyris et al. 2025). The success of these models validates the immense potential of machine learning for complex geoscience problems. However, their limitation is also apparent: they are typically highly optimized for a single objective and often overlook the complex interactions between different physical variables and across multiple scales. This makes them difficult to generalize into a foundation model capable of capturing the full dynamics of the ocean system.
@@
Inspired by the success of foundation models in atmospheric science (Pathak et al. 2022; Bi et al. 2023; Chen et al. 2023a, b; Lam et al. 2023; Oskarsson et al. 2024; Nguyen et al. 2024; Kochkov et al. 2024; Gao et al. 2025), a series of global ocean foundation models have recently emerged, such as AI-GOMS (Xiong et al. 2023), Xihe (Wang et al. 2024c), LangYa (Yang et al. 2024), and WenHai (Cui et al. 2025). These models aim to build a unified digital twin of the ocean and have already surpassed traditional numerical methods in computational efficiency and, in some cases, accuracy (Hao et al. 2025). However, they still face the two core challenges we identified in the introduction. First, they generally adopt standard autoregressive frameworks and have not fundamentally solved the problem of error accumulation in long-term simulations, leading to a decline in forecast fidelity over time. Second, their designs often lack sufficient physical constraints. For example, while models like WenHai (Cui et al. 2025) incorporate domain knowledge into a Swin Transformer (Liu et al. 2021), their feature interaction mechanisms remain relatively physics-agnostic, failing to explicitly model physical processes like gradient changes or variable couplings, nor can they adaptively handle multi-scale dynamics.
@@
$\hat{\mathbf{O}}_{t+1}^{(1)}$ $=\mathcal{M}_{\text{base}}(\mathbf{X}_{t}),$ (2)
@@
We train each component of NeuralOM by minimizing the relative L2-norm between the predicted and ground-truth states. The loss function $\mathcal{L}$ is defined as:
$\mathcal{L}(\Theta)=\mathbb{E}_{t,k}\left[\frac{\left\|\hat{O}_{i,j,k}^{t+1}-O_{i,j,k}^{t+1}\right\|_{2}}{\left\|O_{i,j,k}^{t+1}\right\|_{2}}\right],$ (6)
where $\hat{O}$ and $O$ are the predicted and ground-truth values, respectively, indexed over time $t$, spatial locations $(i,j)$, and variable channels $k$. $\|\cdot\|_{2}$ denotes the Euclidean norm over the spatial grid (i, j) for each channel $k$. ${E}_{t,k}$ denotes the empirical expectation taken over training time steps $t$ and variable channels $k$, i.e., averaging the channel-wise relative errors over training samples.
@@
For long-range simulation and forecasting, we employ a standard autoregressive (or “rollout”) strategy. Starting from an initial state $\mathbf{X}_{0}$, the model continuously feeds its own previous prediction as input for the next step to generate a continuous trajectory spanning several weeks or months. For idealized simulation experiments, we use ground-truth atmospheric data as forcing. This approach eliminates interference from atmospheric forecast errors, allowing for a fairer assessment of the ocean models’ intrinsic performance. For more realistic forecasting experiments, we drive NeuralOM with forecast fields generated by a state-of-the-art atmospheric forecasting model, OneForecast (Gao et al. 2025).
@@
FourCastNet (Pathak et al. 2022) arXiv 2022 5.6295 0.0782 >10>10 0.0190 >10>10 0.0046 >10>10 0.0007
CirT (Liu et al. 2025) ICLR 2025 2.7510 0.0143 2.8686 0.0115 2.9514 0.0099 3.0631 0.0081
WenHai (Cui et al. 2025) Nature Communications 2025 0.7199 0.5065 0.7467 0.4703 0.7700 0.4385 0.8091 0.3851
NeuralOM (ours) 0.6509 0.5297 0.6665 0.5009 0.6795 0.4754 0.7014 0.4314
NeuralOM Improvement 9.59% 4.58% 10.74% 6.49% 11.74% 8.42% 13.32% 12.01%
@@
Our experiments are based on the GLORYS12 global ocean reanalysis dataset, which provides daily mean ocean states since 1993. We split the data into a training set (1993-2017), a validation set (2018-2019), and a test set (2020). The core task is autoregressive forecasting at the Subseasonal-to-Seasonal (S2S) scale, with rollout horizons ranging from several days up to 60 days. The input consists of 93 ocean state variables (including salinity, velocity, temperature at various depths, and sea surface height) and 8 atmospheric variables as external forcing. All spatial data is downsampled to a 24-hour temporal resolution (daily mean for ocean and 12:00 UTC for forcings) and 0.5° spatial resolution. Further details on data processing and variables are available in the Appendix.
@@
We compare NeuralOM against three representative models: (1) FourCastNet (Pathak et al. 2022), a widely recognized atmospheric science foundation model based on Fourier Neural Operators. (2) CirT (Liu et al. 2025), a state-of-the-art atmospheric forecasting model designed for the S2S scale. (3) WenHai (Cui et al. 2025), a state-of-the-art ML-based ocean model that has demonstrated superior performance over traditional numerical models and ML-based ocean foundation models. To ensure a fair comparison, all models, except for the officially released WenHai results, are retrained under our unified framework.
@@
We use four metrics to assess performance: (1) Root Mean Square Error (RMSE) to measure the average prediction error, and (2) Anomaly Correlation Coefficient (ACC) to evaluate the model’s ability to capture deviations from the climatological mean (i.e., anomalies). Additionally, we use (3) Critical Success Index (CSI), and (4) Symmetric Extremal Dependence Index (SEDI) to evaluate the model’s performance on extreme events.
@@
To assess practical applicability, we drive the ocean models with atmospheric forecasts from OneForecast, which introduces real-world uncertainty. As common atmospheric forecast models don’t provide all forcings required by WenHai at the S2S scale, we can’t assess its results. Even in this more challenging scenario, Table 4 shows that NeuralOM retains its lead, achieving significantly better RMSE ($0.68$) and ACC ($0.48$) at a $35$-day lead time. This demonstrates the robustness of NeuralOM for real-world applications, further supported by the high-fidelity visuals in Figure 5(a).
NeuralOM excels at capturing high-impact extreme events.
Variants RMSE ACC
NeuralOM w/o PRC 0.6692 0.5502
NeuralOM w/o PEI 0.6865 0.5166
NeuralOM w/o MANA Sum 0.7037 0.5359
NeuralOM w/o MANA Mean 0.6978 0.5160
NeuralOM 0.6009 0.6058
VARIANTS 23 Layers S 23 Layers T SSH
RMSE ACC RMSE ACC RMSE ACC
NeuralOM w/o Subtract Climatology 0.0515 0.1066 0.1518 0.4532 0.2427 0.5737
NeuralOM w Subtract Climatology 0.0105 0.6412 0.0985 0.6507 0.1375 0.6690
@@
FourCastNet 0.55 0.68 0.68 0.47 1.30 0.19
CirT 1.36 0.14 2.08 0.05 2.61 0.02
Ours 0.52 0.73 0.62 0.59 0.68 0.48
@@
To understand the driving factors behind NeuralOM’s superior performance and to validate the effectiveness of each of our design choices, we conduct detailed ablation studies. We primarily investigate two aspects: the importance of the input representation (i.e., subtracting climatology) and the contribution of each component in our core architecture.
@@
As mentioned in our methodology, we argue that focusing the model on learning “anomalies” rather than the full state is crucial for capturing slow-changing dynamics. To verify this, we train a variant of NeuralOM without subtracting the climatological mean. The results, shown in Table 3, are striking. For sea salinity (S) across 23 depth layers, the model with climatology priors (W Subtract Climatology) achieves an ACC of 0.6412, whereas the model without them (W/O Subtract Climatology) only reaches 0.1066—a performance gap of over 5-fold. This result provides strong evidence that incorporating physical priors like climatology into the input representation is a prerequisite for enabling the model to effectively learn fine-grained physical signals and achieve high-fidelity simulations.
@@
We evaluate the individual contributions of NeuralOM’s core components by removing them one by one. The results, summarized in Table 2, show that removing any single component leads to a significant degradation in performance:
Without Progressive Residual Correction (w/o PRC): Replacing the multi-stage PRC framework with a single model increases RMSE from 0.6009 to 0.6692, confirming that decomposing the complex prediction task into a series of simpler residual correction steps is vital for suppressing error accumulation and ensuring long-term stability.
Without Physics-inspired Edge-Interactions (w/o PEI): Removing PEI, thereby reverting to a traditional, physics-agnostic message passing scheme, caused the ACC to drop from 0.6058 to 0.5166. This confirms that explicitly modeling physical processes like gradients and couplings in message passing improves the model’s physical consistency.
Multi-scale Adaptive Node Aggregation (w/o MANA Sum/Mean): Removing either the sum or the mean aggregation path degrades performance. This confirms the necessity of our adaptive aggregation mechanism, which allows the model to flexibly handle ocean dynamics at different scales.
In summary, these ablation studies confirms that the success of NeuralOM is not coincidental but is the result of a synergistic combination of its carefully designed components from the input representation and macroscopic framework to the microscopic interaction mechanisms.